A boy shoves his stuffed toy zebra down a frictionless chute, starting at a height of 1.21 m above the bottom of the chute and with an initial speed of 1.13 m/s. The toy animal emerges horizontally from the bottom of the chute and continues sliding along a horizontal surface with coefficient of kinetic friction 0.233. How far from the bottom of the chute does the toy zebra come to rest? Take g = 9.81 m/s2.
Answer #1

Conserving energy

Energy at bottom = Energy at top

mgh +0.5mv^2 = E_b

m(9.8*1.21+0.5*1.13^2) = E_b

m*12.5 = E_b

now this energy is used to overcome friction

uN*d = E_b

0.233*m*9.8*d = m*12.5

d = 5.47 m